# ZS4C: Zero-Shot Synthesis of Compilable Code for Incomplete Code Snippets using ChatGPT

### v1

@@ 3.2. Conversational Error Fixing
Through ImportStInfer, we could identify potential missing import statements. However, the code may still not be compiled due to incorrect import statements, bad syntax, typos, and undeclared variables. Therefore we need further validation of our generated code and boost the compilation rate by fixing those errors. Drawing inspiration from recent advancements in the Multi-Agent Conversation and its applications across various tasks (Wu et al., 2023; Zhang et al., 2023; Xi et al., 2023), where multiple LLM-based agents work collaboratively and outperform a single LLM-based agent, we design a ConversationFixing component to further improve the results given by ImportStInfer. Within this component, a Java compiler is integrated into workflow with the LLM (via conversations) to address compilation errors. This compiler not only validates the generated code but also produces an error log that aids in resolving compilation issues. Specifically, we leverage the feedback, i.e., the error log, from the compiler’s validation as context, prompting the LLM to fix the errors for the given code snippet.
@@
Input : Maximum number of interaction between compiler and LLM, $M$
@@ 2 $history$ = ””
3 // Conversation between compiler and LLM up to $K$ times
@@
5       // load dependent libraries dynamically
6       Compiler.loadExternalLibraries($CS$)
7       $CompileResults$ = Compiler.Compile($CS$)
8       // $CS$ is failed to compile
9       if $CompileResults$ is failed then
10             $errorLog$ = getErrorLog(CompilerResults)
@@ 12             $CS$, $history$ = queryLLM($LLM$, $Prompt$)
We present the logic of ConversationFixing in Algorithm 1. The algorithm takes the code snippet with inferred import statements generated by ImportStInfer, denoted as $CS$, the maximum number of interactions between compiler and LLM, represented by $M$, and a large language model, $LLM$, as input, and outputs a most recent fixed code, denoted as $FC$. First, we load external libraries and compile the $CS$ (lines 6-7) to examine if $CS$ can be compiled successfully. If the compilation is successful, we return $CS$ directly. If not, we construct a prompt incorporating the error log and $CS$ (lines 10 - 11). Subsequently, We query the $LLM$ with the constructed prompt, seeking corrections for $CS$ (line 12). This conversation (lines 5 - 15) is repeated up to $M$ times until a successful compilation.
Note that during the compilation, the compiler needs to search for the external libraries to import as external dependencies (line 6). To do so, we dynamically import the external libraries based on the import statements obtained from ImportStInfer or previous round of ConversationFixing. More specially, we construct a knowledge base to store the mappings between fully qualified names (FQN) of classes/functions and their corresponding external libraries (e.g., org.joda.time.Duration $\rightarrow$ joda-time-2.9.9.jar) in an offline manner. See more details of knowledge base construction in Section 4.2. When a piece of code with inferred import statements is input into ConversationFixing, we search the external libraries for each FQN from the import statements and load the corresponding libraries from the constructed knowledge base. It’s worth noting that a single FQN possibly could correspond to multiple external libraries (e.g., different versions of the same libraries provide the same class) (Dong et al., 2022; Subramanian et al., 2014). To mitigate such a case, we selected the one that satisfies pre-defined constraints following the previous study (Subramanian et al., 2014). In our study, we consider the invoked functions and attributes as the constraints. If there exist multiple libraries available after resolving constraints, we select the one that was released most recently. We assume it is realistic for a user to use the most recent libraries. Note that our approach does not guarantee to find all external libraries due to various reasons (e.g., incorrect FQN in import statements), instead our goal is to leverage the message from the compiler to help LLM infer the correct import statements iteratively. For instance, the example in Figure 2 (left), for the import statement “import org.joda.time.PeriodFormatterBuilder;”, no external library was found since the import org.joda.time.PeriodFormatterBuilder (missing format) is incorrect. We still compile the code and the error message generated by the compiler helps LLM fix the error (see more detailed results in Section 5.2).
@@
Presentation_Guide: “Reply with to-the-point answer, no elaboration.” Main_Instruction: “Now fix the error by focusing on fixing the import statements by not using wildcard imports and must not modify code body which means do not change anything inside the class. So, it can be successfully compiled and reply with full code.” # Prompt template for the 1st attempt: {Presentation_Guide} “See the code below:” $\{input\_code\_snippet\_with\_inferred\_import\_statements\}$ “For the above code I got the below error log:” $\{error\_log\}$Main_Instruction} # Prompt starting from the 2nd attempt: $Context$: $\{chat\_history\}$$\{last\_code\_from\_ChatGPT\}$ “You gave the above imports fix in your attempt $\{attempt\_number\}$. But compiler gave this error:” $\{last\_error\_message\}$Main_Instruction}
@@
In each iteration, we update the $CS$ as suggested by $LLM$ and provide the chat history as contextual information. If, after attempting $M$ times, $CS$ remains non-compilable, the latest revised code is returned. Our design mimics how developers use an LLM (e.g., ChatGPT) to fix compilation errors for a piece of code over conversation with LLM (e.g., ChatGPT).
@@ 4.1. Research questions
  • RQ3: Does few-shot learning improve ZS4C?
  • RQ4: How do parameters affect the effectiveness of ZS4C?
4.2. Datasets
@@
We use an established benchmark called StatType-SO, which was used in the related studies (Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2019). The dataset consists of 267 code snippets of six popular Java libraries. Table 1 summarizes the basic statistics of StatType-SO. The six libraries are of different sizes and cover various application domains. This dataset was manually collected from Stack Overflow posts and represents the scenario in which developers use a wide variety of real Java libraries in practice (Phan et al., 2018). This benchmark consists of the code snippets and their dependent libraries. For constructing the knowledge base that stores the mapping between fully qualified names (FQN) of classes/functions and their corresponding external libraries, we downloaded all versions of the most popular 500 libraries from maven repositories [^3] https://mvnrepository.com/popular. We extracted the FQN for all classes/functions from the jar files and created the knowledge base.
Table 1. Basic statistics of StatType-SO.
@@
#Code Snippet #Average LOC #Unique Import Statements
Android 50 38 129
JDK 23 90 103
Hibernate 5 50 123
JodaTime 50 19 48
GWT 50 35 108
XStream 44 37 80
@@ 4.3. Evaluation metrics
Criterion 1: Compilation Rate on the Synthesized Code. To measure the effectiveness of ZS4C in synthesizing compilable code from incomplete code snippets, we use compilation rate, i.e., the ratio of code snippets that successfully compile post-repair to the total number of snippets evaluated, following previous studies (Dong et al., 2022; Terragni et al., 2016).
@@ 4.4. Base large language model
In this study, we choose ChatGPT (OpenAI, 2023a) as our base LLM. We select ChatGPT as it is recognized to be one of the most powerful LLMs and has gained wide attention in recent months by demonstrating the outstanding capability of solving tasks in various areas (Frieder et al., 2023), including software engineering (Li et al., 2023; Surameery and Shakor, 2023; Ma et al., 2023; Xie et al., 2023; Xia and Zhang, 2023). We accessed ChatGPT, i.e., the default GPT-3.5-Turbo model, via the OpenAI API (OpenAI, 2023b).
@@ 4.6.1. Approach of RQ1
  • SnR: The state-of-the-art technique for synthesizing incomplete code snippets by inferring missing import statements and making them compilable (Dong et al., 2022). SnR has been demonstrated to archive better performance than other approaches (e.g., COSTER (Saifullah et al., 2019), Baker (Subramanian et al., 2014), and StatType (Phan et al., 2018)) on StatType-SO. Therefore, in this study, we only compare with SnR. We leveraged the replication package of SnR as provided in their paper. [^4] We found disparity between the number reported in the paper and the number derived from their replication package. We present the numbers derived from their replication package in this paper. We do not compare with CSNIPPEx because this Eclipse plugin is no longer accessible.
  • BasePrompt: We also consider a baseline that employs a straightforward prompt to instruct a LLM to make the given code snippet compilable. Specifically, we replace the Main_Instruction as “Make the code below compilable: {$input\_code\_snippet$}” in Prompt Template-1 (ref. Section 3.1) and while retaining the remainder of the prompt template unchanged.
We evaluate the effectiveness of ZS4C against the two baselines on the two criteria outlined in Section 4.3. To ensure the reliability of our results and mitigate potential biases arising from the inherent randomness of the LLM, we run both ZS4C and BasePrompt five times and take the median value as the final result.
@@ 4.6.2. Approach of RQ2
[…]
@@ 4.6.3. Approach of RQ3
Selecting good demonstrations is essential for few-shot in-context learning. Prior studies show that selecting demonstration examples with higher similarity (to the input) can improve in-context learning performance (Nashid et al., 2023; Zhou et al., 2022a). Thus, in this RQ, we incorporate code snippets that are most similar to each input code snippet, along with their corresponding import statements, as demonstrations within the prompt. Specifically, we employ the commonly used CodeBERT (Feng et al., 2020) model that was trained for code-related tasks to learn a vector representation for each code snippet in StatType-SO. Based on the code representations output by CodeBERT, we select the code snippets presenting the highest cosine similarities with the input code snippet. Due to the window size token limitation of ChatGPT (i.e., 4k), we only put the most similar code snippet and its import statements in the prompt as the demonstration example. We put the retrieved demonstrated example after the Main_Instruction in Prompt Template - 1 as shown in Section 3.1.1. In other words, we create a one-shot version of ZS4C.
4.6.4. Approach of RQ4
ZS4C has two parameters: $K$ the number of sampling in self-consistency, and $M$ the maximum rounds of conversation between the compiler and the LLM in Conversational Error Fixing. In this RQ, we investigate the impact of each parameter on the effectiveness of ZS4C, where we set $K$ to 10, 20, and 30 and set $M$ to 5, 10, and 15, respectively.

### v3

@@ 3.2. Conversational Error Fixing
Through ImportStInfer, we could identify potential missing import statements. However, the code may still not be compiled due to incorrect import statements, bad syntax, typos, and undeclared variables. Therefore we need further validation of our generated code and boost the compilation rate by fixing those errors. Drawing inspiration from recent advancements in the Multi-Agent Conversation and its applications across various tasks (Wu et al., 2023; Zhang et al., 2023; Xi et al., 2023), where multiple LLM-based agents work collaboratively and outperform a single LLM-based agent, we design a ConversationFixing component to further improve the results given by ImportStInfer. Within this component, a validator (e.g., a compiler for Java code) is integrated into the workflow with the LLM (via conversations) to address compilation errors. This validator not only validates the generated code but also produces an error log that aids in resolving compilation issues. Specifically, we leverage the feedback, i.e., the error log, from the validator as context, prompting the LLM to fix the errors for the given code snippet.
1
@@
Input : Maximum number of interaction between validator and LLM, $M$
@@ 3 $history$ = ””
4 // Conversation between validator and LLM up to $K$ times
@@
6       $ValidationResults$ = Validator.validate($CS$)
7       // $CS$ failed to pass the validator
8       if $ValidationResults$ is failed then
9             $errorLog$ = getErrorLog($ValidationResults$)
@@ 11             $CS$, $history$ = queryLLM($LLM$, $Prompt$)
We present the logic of ConversationFixing in Algorithm 1. The algorithm takes the code snippet with inferred import statements generated by ImportStInfer, denoted as $CS$, the maximum number of interactions between validator and LLM, represented by $M$, and a large language model, $LLM$, as input, and outputs a most recent fixed code, denoted as $FC$. First, we examine if $CS$ can pass the validator successfully (line 5). If the validation is successful, we return $CS$ directly. If not, we construct a prompt incorporating the error log and $CS$ (lines 8 - 9). Subsequently, We query the $LLM$ with the constructed prompt, seeking corrections for $CS$ (line 10). This conversation (lines 4 - 14) is repeated up to $M$ times until a successful compilation.
Prompt Template - 2 for Conversational Error Fixing Presentation_Guide: “Reply with to-the-point answer, no elaboration.”
@@
Main_Instruction (for Java): “Now fix the error by focusing on fixing the import statements by not using wildcard imports and must not modify code body which means do not change anything inside the class. So, it can be successfully compiled and reply with full code.”
Main_Instruction (for Python): “Now fix the error by focusing on fixing the import statements. So, it can be run successfully and reply with full code.”
# Prompt template for the 1st attempt:
{Presentation_Guide}
“See the code below:”
$\{input\_code\_snippet\_with\_inferred\_import\_statements\}$
“For the above code I got the below error log:”
$\{error\_log\}$
{Main_Instruction}
# Prompt starting from the 2nd attempt:
$Context$:
$\{chat\_history\}$
$\{last\_code\_from\_ChatGPT\}$
“You gave the above imports fix in your attempt $\{attempt\_number\}$. But compiler gave this error:”
$\{last\_error\_message\}$
{Main_Instruction}
@@
In each iteration, we update the $CS$ as suggested by $LLM$ and provide the chat history as contextual information. If, after attempting $M$ times, $CS$ remains unsuccessful, the latest revised code is returned. Our design mimics how developers use an LLM (e.g., ChatGPT) to fix compilation errors for a piece of code over conversation with LLM (e.g., ChatGPT).
@@
Note that during the validation, if a compiler is used as a validator for programming languages such as Java and C++ which needs compilation first, the compiler needs to search for the external libraries to import as external dependencies. To do so, we dynamically import the external libraries based on the import statements obtained from ImportStInfer or previous round of ConversationFixing. More specially, we construct a knowledge base to store the mappings between fully qualified names (FQN) of classes/functions and their corresponding external libraries (e.g., org.joda.time.Duration $\rightarrow$ joda-time-2.9.9.jar) in an offline manner. See more details of knowledge base construction in Section 4.2.1. When a piece of code with inferred import statements is input into ConversationFixing, we search the external libraries for each FQN from the import statements and load the corresponding libraries from the constructed knowledge base. It’s worth noting that a single FQN possibly could correspond to multiple external libraries (e.g., different versions of the same libraries provide the same class) (Dong et al., 2022; Subramanian et al., 2014). To mitigate such a case, we selected the one that satisfies pre-defined constraints following the previous study (Subramanian et al., 2014). In our study, we consider the invoked functions and attributes as the constraints. If there exist multiple libraries available after resolving constraints, we select the one that was released most recently. We assume it is realistic for a user to use the most recent libraries. Note that our approach does not guarantee to find all external libraries due to various reasons (e.g., incorrect FQN in import statements), instead our goal is to leverage the message from the compiler to help LLM infer the correct import statements iteratively. For instance, the example in Figure 2 (left), for the import statement “import org.joda.time.PeriodFormatterBuilder;”, no external library was found since the import org.joda.time.PeriodFormatterBuilder (missing format) is incorrect. We still compile the code and the error message generated by the compiler helps LLM fix the error (see more detailed results in Section 5.2).
For Java code snippets, we select the standard compiler Conversational Error Fixing component. For Python code snippets, we use Pyflakes (Developers, 2024), which is a static analysis tool for Python that checks for errors (syntax errors, undefined variables, and other common issues) in Python source files. For this study, we utilized Pyflakes version 3.2.0 to analyze our codebase before executing it with a Python interpreter.
@@ 4.1. Research questions
  • RQ3: How do parameters affect the effectiveness of ZS4C?
  • RQ4: Does few-shot learning improve ZS4C?
4.2. Data preparation
4.2.1. Datasets
@@
To evaluate the effectiveness of ZS4C, we first use an established benchmark called StatType-SO, which was used in the related studies (Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2019). The dataset consists of 267 code snippets of six popular Java libraries. The six libraries are of different sizes and cover various application domains. This dataset was manually collected from Stack Overflow posts and represents the scenario in which developers use a wide variety of real Java libraries in practice (Phan et al., 2018). This benchmark consists of the code snippets and their dependent libraries.
To complement and diversify StatType-SO, which only consists of Java libraries, we also constructed a Python dataset following previous studies (Phan et al., 2018; Subramanian et al., 2014). We first selected the top 20 most popular Python libraries (Team, 2024). For each library $L$ of interest, we searched for all the posts from the past 10 years that use $L$ via Stack Overflow tags on StackExchange Data Explorer [^3] https://data.stackexchange.com/ and downloaded them. To ensure the quality of the posts, we only included those with a score greater than 5. We considered code snippets that have more than ten lines of code in both questions and answers. We ended up with 539 code snippets across the 20 studied libraries.
To construct the benchmark (i.e., the correct import statements for each Python code snippet), the first author manually collected the required modules by examining the posts on Stack Overflow that contain the corresponding code snippet and searching on Google. Note that some code snippets initially came with partially imported modules; we removed those modules when constructing the dataset to avoid data leakage.
Table 1 summarizes the basic statistics of StatType-SO and Python-SO.
4.2.2. Knowledge base construction
In Section 3.2, we use a compiler to validate the code snippet and provide feedback (i.e., error log) to the LLM for error fixing for Java code snippets. To make the compiler ready for compilation, we need to import proper external libraries for dependencies. For this, we need to construct the knowledge base that stores the mapping between fully qualified names (FQN) of classes/functions and their corresponding external libraries. To do so, we downloaded all versions of the most popular 500 libraries from maven repositories [^4] https://mvnrepository.com/popular. We then performed the following steps to extract the FQN and the last modification date for each external library.
  • FQN Extraction: We first decompress the jar file for each external library. Next, we use the javap tool [^5] https://docs.oracle.com/javase/8/docs/technotes/tools/windows/javap.html, which is a Java class file disassembler, to analyze the binary class files extracted from the JAR. The output from javap is parsed to identify all non-abstract, non-interface class definitions, and then extract their fully qualified names.
  • Lastest Modification Date: Note that we also need the modified date of each library to select the most recent library if multiple libraries meet constraints. To retrieve the last modification date of the JAR file, we examine the metadata associated with the file within the JAR. This date provides a timestamp indicating the most recent update or release of the JAR file. Note that for some jars we were not able to extract the valid modification date. Therefore, we added them manually from the Maven repository.
We then create the inverse index for each pair of FQN and its corresponding library (e.g., org.joda.time.Duration $\rightarrow$ joda-time-2.9.9.jar).
For Python code snippets, we do not use a compiler, since it is an interpreted language. Therefore, we did not construct a separate knowledge base. We use Pyflakes (Developers, 2024) to perform the validation and provide feedback in Section 3.2. After that, we execute Python code by invoking the Python interpreter as a subprocess within our scripts. To ensure a consistent and isolated environment for running the Python code, we use a virtual environment. We identified and listed the necessary libraries for Python-SO, ensuring their installation in the latest versions within this virtual environment. This strategy enabled us to manage dependencies efficiently and avoid conflicts with other projects.
Table 1. Basic statistics of StatType-SO and Python-SO, including the number of code snippets, the average LOC, and unique import statement in each library. Note that #Unique Import Statements refers to the total number of unique import statements that need to be predicted correctly for each library.
StatType-SO
@@ 4.2.2. Knowledge base construction
Library #Code Snippet Mean #LOC Max #LOC Standard Deviation of #LOC #Unique Import Statements
Android 50 35 192 34 129
JDK 23 82 300 64 103
Hibernate 50 42 136 29 123
JodaTime 50 17 49 7 48
GWT 50 31 111 16 108
XStream 44 35 100 18 80
Python-SO
Library #Code Snippet Mean #LOC Max #LOC Standard Deviation of #LOC #Unique Import Statements
BeautifulSoup 35 13 29 6 31
Flask 23 14 29 6 32
Gensim 11 15 28 5 16
Keras 17 24 116 24 38
Lightgbm 7 18 36 11 13
Matplotlib 49 13 36 7 18
NLTK 18 14 46 9 26
Numpy 17 22 109 24 21
OpenCV 50 19 109 16 14
Pandas 8 18 36 10 9
Plotly 53 20 76 11 27
Pytorch 26 16 30 7 24
Scikit-learn 17 15 29 6 24
Scipy 37 16 65 10 44
Scrapy 49 19 113 18 74
Seaborn 48 15 53 10 17
Spacy 26 14 26 4 26
Tensorflow 16 36 132 36 23
Theano 11 11 18 3 20
XGBoost 21 21 75 14 35
@@ 4.3. Evaluation metrics
Criterion 1: Compilation Rate on the Synthesized Code. To measure the effectiveness of ZS4C in synthesizing compilable code from incomplete code snippets, we use compilation rate (CR), i.e., the ratio of code snippets that successfully compile post-repair to the total number of snippets evaluated, following previous studies (Dong et al., 2022; Terragni et al., 2016). Note that Python code does not need to be compiled, therefore, we use a standard Python Interpreter to execute the code, and if it passes, it typically refers to the successful execution. For simplicity, throughout the paper, we consider it compilable and refer to it as the compilation rate (CR).
@@ 4.4. Base large language model
In this study, we choose GPT-3.5 and GPT-4 as our base LLMs (OpenAI, 2023a). We select them as they are recognized to be the most powerful LLMs and have gained wide attention in recent months by demonstrating the outstanding capability of solving tasks in various areas (Frieder et al., 2023), including software engineering (Li et al., 2023; Surameery and Shakor, 2023; Ma et al., 2023; Xie et al., 2023; Xia and Zhang, 2023). We accessed the “gpt-3.5-turbo-0125” for GPT-3.5 and “gpt-4o” model for GPT-4, via the OpenAI API (OpenAI, 2023b, c). We denote the ZS4C with GPT-3.5 and GPT-4 as ZS4CGPT-3.5 and ZS4CGPT-4, respectively.
@@ 4.6.1. Approach of RQ1
  • SnR: The state-of-the-art technique for synthesizing incomplete code snippets by inferring missing import statements and making them compilable (Dong et al., 2022). SnR has been demonstrated to archive better performance than other approaches (e.g., COSTER (Saifullah et al., 2019), Baker (Subramanian et al., 2014), and StatType (Phan et al., 2018)) on StatType-SO. Therefore, in this study, we only compare with SnR. We leveraged the replication package of SnR as provided in their paper. [^6] We found the disparity between the number reported in the paper and the number derived from their replication package. We present the numbers derived from their replication package in this paper. We do not compare with CSNIPPEx because this Eclipse plugin is no longer accessible.
  • BasePrompt: We also consider a baseline that employs a straightforward prompt to instruct an LLM to make the given code snippet compilable. Specifically, we replace the Main_Instruction as “Make the code below compilable: {$input\_code\_snippet$}” in Prompt Template-1 (ref. Section 3.1) and while retaining the remainder of the prompt template unchanged. We denote the BasePrompt with GPT-3.5 and GPT-4 as BasePromptGPT-3.5 and BasePromptGPT-4, respectively.
We evaluate the effectiveness of ZS4C with GPT-3.5 and GPT4 against the two baselines on the two criteria outlined in Section 4.3 on StatType-SO. Since SnR is designed for Java and is not appliable to Python code, we focus on comparing ZS4C with BasePrompt for Python-SO dataset. To ensure the reliability of our results and mitigate potential biases arising from the inherent randomness of the LLM, we run both ZS4C and BasePrompt five times and take the median value as the final result.
@@ 4.6.2. Approach of RQ2
In this RQ, we carry out an ablation study to understand the contribution of each component (i.e., ConversationFixing and ImportStInfer) in ZS4C. Note that for ConversationFixing, the input code snippet will be the original code snippet without any inferred import statements. We are also interested in understanding the contribution of the self-consistency method using in ImportStInfer. Therefore, we remove self-consistency from ImportStInfer (refer as to ImportStInfernoSC) and compare it with ImportStInfer. We then evaluate the effectiveness of the above three approaches following the same methodology in RQ1, and compare them with ZS4C. Note that using the API of GPT-4 is expensive, and we observe similar performance between GPT-4 and GPT-3.5. Therefore, we carry out the ablation study with the more economical model GPT-3.5.
@@ 4.6.3. Approach of RQ3
ZS4C has two parameters: $K$ the number of sampling in self-consistency, and $M$ the maximum rounds of conversation between the compiler and the LLM in Conversational Error Fixing. In this RQ, we investigate the impact of each parameter on the effectiveness of ZS4C, where we set $K$ to 10, 20, and 30 and set $M$ to 5, 10, and 15, respectively. Similar to RQ2, we carry out the ablation study with GPT-3.5.
4.6.4. Approach of RQ4
@@
Selecting good demonstrations is essential for few-shot in-context learning. Prior studies show that selecting demonstration examples with higher similarity (to the input) can improve in-context learning performance (Nashid et al., 2023; Zhou et al., 2022a). Thus, in this RQ, we incorporate code snippets that are most similar to each input code snippet, along with their corresponding import statements, as demonstrations within the prompt. Specifically, we employ the commonly used CodeBERT (Feng et al., 2020) model that was trained for code-related tasks to learn a vector representation for each code snippet in StatType-SO. Based on the code representations output by CodeBERT, we select the code snippets presenting the highest cosine similarities with the input code snippet. When testing the performance of ZS4C with few-shot in-context learning, we selected different configurations based on the capabilities of the language models and their token window size limitations. Table 2 presents the five-number summary of the prompt token size for different settings. As we can see the median size for three-shot prompts is 7,100, which indicates that the majority of the prompts exceed the window size of GPT-3.5. Therefore, for GPT-3.5, we utilized a one-shot learning approach, while for GPT-4, we employed both one-shot and three-shot learning approaches. This decision was made due to the token window size constraints (i.e., 4k input tokens for GPT-3.5 and 128k input tokens for GPT-4). We put the retrieved demonstrated example after the Main_Instruction in Prompt Template - 1 as shown in Section 3.1.1.
Table 2. Summary of prompt size for different settings in terms of input token size distributions, including the five-number summary statistics (Min: Minimum, Q1: 25th percentile, Median: 50th percentile, Q3: 75th percentile, Max: Maximum).
StatType-SO Python-SO
Min Q1 Median Q3 Max Min Q1 Median Q3 Max
Three-shot 4,920 5,654 7,100 10,380 100,650 2,370 4,200 5,790 8,430 42,060
One-shot 1,200 1,934 3,320 4,840 65,020 1,580 2,800 3,860 5,620 28,040
Zero-shot 620 1,082 1658 2,440 32,530 790 1,400 1,930 2,810 14,020